OpenStudy (anonymous):

Simplify (s)(-3st)(-).

OpenStudy (danjs):

what is (-)

OpenStudy (anonymous):

ops its \[\frac{ 1 }{ 3}\]

OpenStudy (danjs):

since it is all multiplication, you can combine them all into one term

OpenStudy (anonymous):

yea but what do i do with the letters

OpenStudy (danjs):

-3*s*s*t*(1/3)

OpenStudy (danjs):

remember... \[s*s = s^{(1+1)} = s^2\]

OpenStudy (danjs):

-3/3 = -1

OpenStudy (anonymous):

so 64? and -1

OpenStudy (danjs):

hmm.. here

OpenStudy (danjs):

\[\frac{ -3*s*s*t }{ 3 } = \frac{ -3 }{ 3 }*s^2*t\]

OpenStudy (danjs):

where did you see 64

OpenStudy (anonymous):

i did the 8 example thing you showed me

OpenStudy (danjs):

s*s = s^2 that is all you can do... the numbers and letters are not able to combine

OpenStudy (danjs):

\[\frac{ -3*s*s*t }{ 3 } = \frac{ -3 }{ 3 }*s^2*t = (-1)*s^2*t\] that is all you can do

OpenStudy (anonymous):

so -3 1/3 t?

OpenStudy (usukidoll):

the 3's cancel .... as what @DanJS typed earlier.

OpenStudy (danjs):

-s^2*t

OpenStudy (anonymous):

@DanJS i was typing that

OpenStudy (danjs):

you can combine only the like quantities... numbers or variables

OpenStudy (danjs):

cool.. goodluck

OpenStudy (anonymous):

thanks for your help